OpenStudy (hang254):

Find the limit of the function algebraically. lim x goes to 0 (x^2+2x)/(x^4)

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{x^2+2x}{x^4}\] You could factor out an x on top and and x on bottom

OpenStudy (hang254):

so \[\lim (x+2)/x^3\]

myininaya (myininaya):

what happens when you plug in 0

OpenStudy (hang254):

you get 0

myininaya (myininaya):

0+2 isn't 0

OpenStudy (hang254):

2/0 = 0

myininaya (myininaya):

2/0 is not 0

myininaya (myininaya):

the only thing that can possibly make a fraction 0 is the top

myininaya (myininaya):

if the top is not 0, then you definitely know the fraction is never 0

OpenStudy (hang254):

undefined

OpenStudy (hang254):

oh

myininaya (myininaya):

and if you have 0/0 then that would be indeterminate form but we have 2/0 which is undefined or does not exist

myininaya (myininaya):

0/0 means you have to do more work 2/0 means you are done 0/2 would be 0

OpenStudy (hang254):

alright, thanks for explaining that to me

myininaya (myininaya):

np